During the test, (C10H22) is burnt using an air-fuel ratio of 13:1 by weight. Determine the
complete volumetric analysis of the products of combustion, if the whole amount of
hydrogen burns to form water vapour and there is neither any free oxygen nor any free
carbon. The carbon burns to CO2 and CO.
Answer #1

Solution [Actual combustion equation] Ciotz2+ a[02+3.56N2] b Co₂ tcco + d H₂O + enz. [to find a, b, c, d & e] a [ 32+ 3.76X282b + 6 + 11 = 2 x 13.447 i, 2b+c= 15.834 from 0 & 2 b=5.894 4.106 By n balance ee= 2x ax 3.76 :. e - 3.76X 13.447 e=50-56 V

Please Rate,

Thank You!!!!